OpenStudy (anonymous):

John is 3 times as old as his brother Sam. In 5 years, John will be twice as old as Sam will be then. Find their present ages.

OpenStudy (anonymous):

j=3s j+5=2(s+5) putting value of j; 3s+5=2s+10 3s-2s=10-5 s=5 now as j=3s so j=3(5)=15